OpenStudy (anonymous):

How many real number solutions does this equation have? –7x^2 + 6x + 3 = 0

OpenStudy (inkyvoyd):

2

OpenStudy (anonymous):

Use the abc-formula to get the Discriminant. Because: Discriminant < 0 = no answers for the equation Discriminant = 0 : 1 answer to the equation Discriminant > 0 : 2 answers to the equation Fill in the a, b and c: a = -7 b = 6 c = 3 Get the discriminant: \[Discrimant = b^2-4ac\] so: \[6^2-4\times(-7) \times 3 = 36 - (-84) = 36 +84 = 120\] 120 > 0 : so there are two answers!